OpenStudy (anonymous):

An uncovered human skeleton is discovered to only have 1/15 of the original amount of Carbon 14 for a human. How long has the person been dead? (Carbon 14 half-life is 5568 years)

OpenStudy (anonymous):

@Sourav12084

OpenStudy (amistre64):

full amount / 2^(hl) = current amount ... sounds familiar

OpenStudy (amistre64):

1/15 is about equal to 1/2^k where k = halflife/#periods dead

OpenStudy (anonymous):

So, it would be FA = 2^(5568) = 1/15 right?

OpenStudy (amistre64):

spose the half life is 3 years; and there is 1/20 remaining 20 = 2^k when k = ln(20)/ln(2) k = 3/n n = 3/k

OpenStudy (amistre64):

\[\frac{FA}{2^{h/n}}=CA\] \[FA=CA*2^{h/n}\] \[\frac{FA}{CA}=2^{h/n}\] goes something along those lines i believe

OpenStudy (amistre64):

2^(5568/n) = 1/15

OpenStudy (anonymous):

\[(1/2)^{n}=1/2^{15}\], find out n....then multiply n with 5568 years.... n is the no of half lives....so (1/2)^n means the amount of material left after n half lives... if you find out n, then just multiply it by 5568 years...

OpenStudy (anonymous):

So, n = 15? 15*5568 = 83520

OpenStudy (amistre64):

one more attempt to organize my thoughts lol at 0 periods there is 1/2^0 total remaining at 1 periods there is 1/2^1 total remaining at 2 periods there is 1/2^2 total remaining at 3 periods there is 1/2^3 total remaining ... at n periods there is 1/2^n total remaining 1/2^n = 1/15, 2^n = 15 when n=log(15)/log(2) n*periods is the number of periods covered

OpenStudy (anonymous):

i'm a little confused :P I feel like I'm trying to solve 2 different problems

OpenStudy (amistre64):

in a sense you are :) the key is in determining the number of unit periods, then multiplying that by the period amount

OpenStudy (anonymous):

so was the "n = 15? 15*5568 = 83520" part correct?

OpenStudy (amistre64):

not that i see ....

OpenStudy (amistre64):

2^n = 15 . not n=15

OpenStudy (anonymous):

then after that, how would i proceed?

OpenStudy (amistre64):

n*(period) of course ...

OpenStudy (anonymous):

that would make it 15*83520 (hl*2). correct?

OpenStudy (amistre64):

n*(5568) but the solution to n is determined by 2^n = 15 what is n?

OpenStudy (anonymous):

2^n = 15 -> n = 15 15*5568 = 83520 i think i might be wrong somewhere

OpenStudy (amistre64):

2^15 does not equal 15

OpenStudy (amistre64):

\[2^n = 15\] \[log(2^n) = log(15)\] \[n~log(2) = log(15)\] \[n =\frac{ log(15)}{log(2)}\] n*5568 will be how long its been dead.

OpenStudy (anonymous):

i'm confused then. i'm not sure how to get n out of it. :P

OpenStudy (anonymous):

so that would be 21753.5668 years? :/

OpenStudy (amistre64):

that does seem like a better number yes :)

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

youre welcome